\begin{thebibliography}{35}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bast and Celikik(2014)]{bast2014efficient}
Hannah Bast and Marjan Celikik.
\newblock Efficient index-based snippet generation.
\newblock \emph{ACM Transactions on Information Systems (TOIS)}, 32\penalty0 (2):\penalty0 1--24, 2014.

\bibitem[Bender et~al.(2021)Bender, Gebru, McMillan-Major, and Shmitchell]{bender2021dangers}
Emily~M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell.
\newblock On the dangers of stochastic parrots: Can language models be too big?
\newblock In \emph{Proceedings of the 2021 ACM conference on fairness, accountability, and transparency}, pages 610--623, 2021.

\bibitem[Chen et~al.(2020)Chen, Syed, Stein, Hagen, and Potthast]{chen2020abstractive}
Wei-Fan Chen, Shahbaz Syed, Benno Stein, Matthias Hagen, and Martin Potthast.
\newblock Abstractive snippet generation.
\newblock In \emph{Proceedings of The Web Conference 2020}, pages 1309--1319, 2020.

\bibitem[Chung et~al.(2022)Chung, Hou, Longpre, Zoph, Tay, Fedus, Li, Wang, Dehghani, Brahma, et~al.]{chung2022scaling}
Hyung~Won Chung, Le~Hou, Shayne Longpre, Barret Zoph, Yi~Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, et~al.
\newblock Scaling instruction-finetuned language models.
\newblock \emph{arXiv preprint arXiv:2210.11416}, 2022.

\bibitem[Feng et~al.(2023)Feng, Balachandran, Bai, and Tsvetkov]{feng2023factkb}
Shangbin Feng, Vidhisha Balachandran, Yuyang Bai, and Yulia Tsvetkov.
\newblock Factkb: Generalizable factuality evaluation using language models enhanced with factual knowledge.
\newblock \emph{arXiv preprint arXiv:2305.08281}, 2023.

\bibitem[Gao et~al.(2024)Gao, Xiong, Gao, Jia, Pan, Bi, Dai, Sun, Wang, and Wang]{gao2024retrievalaugmentedgenerationlargelanguage}
Yunfan Gao, Yun Xiong, Xinyu Gao, Kangxiang Jia, Jinliu Pan, Yuxi Bi, Yi~Dai, Jiawei Sun, Meng Wang, and Haofen Wang.
\newblock Retrieval-augmented generation for large language models: A survey, 2024.
\newblock URL \url{https://arxiv.org/abs/2312.10997}.

\bibitem[Holtzman et~al.(2019)Holtzman, Buys, Du, Forbes, and Choi]{holtzman2019curious}
Ari Holtzman, Jan Buys, Li~Du, Maxwell Forbes, and Yejin Choi.
\newblock The curious case of neural text degeneration.
\newblock \emph{arXiv preprint arXiv:1904.09751}, 2019.

\bibitem[Ji et~al.(2023)Ji, Lee, Frieske, Yu, Su, Xu, Ishii, Bang, Madotto, and Fung]{ji2023survey}
Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye~Jin Bang, Andrea Madotto, and Pascale Fung.
\newblock Survey of hallucination in natural language generation.
\newblock \emph{ACM Computing Surveys}, 55\penalty0 (12):\penalty0 1--38, 2023.

\bibitem[Jiang et~al.(2023)Jiang, Sablayrolles, Mensch, Bamford, Chaplot, Casas, Bressand, Lengyel, Lample, Saulnier, et~al.]{jiang2023mistral}
Albert~Q Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra~Singh Chaplot, Diego de~las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, et~al.
\newblock Mistral 7b.
\newblock \emph{arXiv preprint arXiv:2310.06825}, 2023.

\bibitem[Jin et~al.(2019)Jin, Dhingra, Liu, Cohen, and Lu]{jin-etal-2019-pubmedqa}
Qiao Jin, Bhuwan Dhingra, Zhengping Liu, William Cohen, and Xinghua Lu.
\newblock {P}ub{M}ed{QA}: A dataset for biomedical research question answering.
\newblock In Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan, editors, \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 2567--2577, Hong Kong, China, November 2019. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/D19-1259}.
\newblock URL \url{https://aclanthology.org/D19-1259}.

\bibitem[Kaplan et~al.(2020)Kaplan, McCandlish, Henighan, Brown, Chess, Child, Gray, Radford, Wu, and Amodei]{kaplan2020scaling}
Jared Kaplan, Sam McCandlish, Tom Henighan, Tom~B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei.
\newblock Scaling laws for neural language models.
\newblock \emph{arXiv preprint arXiv:2001.08361}, 2020.

\bibitem[Lewis et~al.(2020)Lewis, Perez, Piktus, Petroni, Karpukhin, Goyal, K{\"u}ttler, Lewis, Yih, Rockt{\"a}schel, et~al.]{lewis2020retrieval}
Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich K{\"u}ttler, Mike Lewis, Wen-tau Yih, Tim Rockt{\"a}schel, et~al.
\newblock Retrieval-augmented generation for knowledge-intensive nlp tasks.
\newblock \emph{Advances in Neural Information Processing Systems}, 33:\penalty0 9459--9474, 2020.

\bibitem[Li et~al.(2023)Li, Holtzman, Fried, Liang, Eisner, Hashimoto, Zettlemoyer, and Lewis]{li-etal-2023-contrastive}
Xiang~Lisa Li, Ari Holtzman, Daniel Fried, Percy Liang, Jason Eisner, Tatsunori Hashimoto, Luke Zettlemoyer, and Mike Lewis.
\newblock Contrastive decoding: Open-ended text generation as optimization.
\newblock In Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki, editors, \emph{Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 12286--12312, Toronto, Canada, July 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.acl-long.687}.
\newblock URL \url{https://aclanthology.org/2023.acl-long.687}.

\bibitem[Lin(2004)]{lin-2004-rouge}
Chin-Yew Lin.
\newblock {ROUGE}: A package for automatic evaluation of summaries.
\newblock In \emph{Text Summarization Branches Out}, pages 74--81, Barcelona, Spain, July 2004. Association for Computational Linguistics.
\newblock URL \url{https://aclanthology.org/W04-1013}.

\bibitem[Liu et~al.(2021)Liu, Sap, Lu, Swayamdipta, Bhagavatula, Smith, and Choi]{liu-etal-2021-dexperts}
Alisa Liu, Maarten Sap, Ximing Lu, Swabha Swayamdipta, Chandra Bhagavatula, Noah~A. Smith, and Yejin Choi.
\newblock {DE}xperts: Decoding-time controlled text generation with experts and anti-experts.
\newblock In Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli, editors, \emph{Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers)}, pages 6691--6706, Online, August 2021. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2021.acl-long.522}.
\newblock URL \url{https://aclanthology.org/2021.acl-long.522}.

\bibitem[Lu et~al.(2021)Lu, West, Zellers, Le~Bras, Bhagavatula, and Choi]{lu-etal-2021-neurologic}
Ximing Lu, Peter West, Rowan Zellers, Ronan Le~Bras, Chandra Bhagavatula, and Yejin Choi.
\newblock {N}euro{L}ogic decoding: (un)supervised neural text generation with predicate logic constraints.
\newblock In Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz~Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou, editors, \emph{Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pages 4288--4299, Online, June 2021. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2021.naacl-main.339}.
\newblock URL \url{https://aclanthology.org/2021.naacl-main.339}.

\bibitem[Marchionini(2006)]{marchionini2006exploratory}
Gary Marchionini.
\newblock Exploratory search: from finding to understanding.
\newblock \emph{Communications of the ACM}, 49\penalty0 (4):\penalty0 41--46, 2006.

\bibitem[Metzler et~al.(2021)Metzler, Tay, Bahri, and Najork]{metzler2021rethinking}
Donald Metzler, Yi~Tay, Dara Bahri, and Marc Najork.
\newblock Rethinking search: making domain experts out of dilettantes.
\newblock In \emph{Acm sigir forum}, volume~55, pages 1--27. ACM New York, NY, USA, 2021.

\bibitem[Narayan et~al.(2018)Narayan, Cohen, and Lapata]{narayan-etal-2018-dont}
Shashi Narayan, Shay~B. Cohen, and Mirella Lapata.
\newblock Don{'}t give me the details, just the summary! topic-aware convolutional neural networks for extreme summarization.
\newblock In Ellen Riloff, David Chiang, Julia Hockenmaier, and Jun{'}ichi Tsujii, editors, \emph{Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing}, pages 1797--1807, Brussels, Belgium, October-November 2018. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/D18-1206}.
\newblock URL \url{https://aclanthology.org/D18-1206}.

\bibitem[Nema et~al.(2017)Nema, Khapra, Laha, and Ravindran]{nema-etal-2017-diversity}
Preksha Nema, Mitesh~M. Khapra, Anirban Laha, and Balaraman Ravindran.
\newblock Diversity driven attention model for query-based abstractive summarization.
\newblock In Regina Barzilay and Min-Yen Kan, editors, \emph{Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1063--1072, Vancouver, Canada, July 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1098}.
\newblock URL \url{https://aclanthology.org/P17-1098}.

\bibitem[O'Brien and Lewis(2023)]{o2023contrastive}
Sean O'Brien and Mike Lewis.
\newblock Contrastive decoding improves reasoning in large language models.
\newblock \emph{arXiv preprint arXiv:2309.09117}, 2023.

\bibitem[Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu]{raffel2020exploring}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J Liu.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{The Journal of Machine Learning Research}, 21\penalty0 (1):\penalty0 5485--5551, 2020.

\bibitem[See et~al.(2017)See, Liu, and Manning]{see-etal-2017-get}
Abigail See, Peter~J. Liu, and Christopher~D. Manning.
\newblock Get to the point: Summarization with pointer-generator networks.
\newblock In Regina Barzilay and Min-Yen Kan, editors, \emph{Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1073--1083, Vancouver, Canada, July 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1099}.
\newblock URL \url{https://aclanthology.org/P17-1099}.

\bibitem[Shah and Bender(2022)]{shah2022situating}
Chirag Shah and Emily~M Bender.
\newblock Situating search.
\newblock In \emph{Proceedings of the 2022 Conference on Human Information Interaction and Retrieval}, pages 221--232, 2022.

\bibitem[Shi et~al.(2023)Shi, Han, Lewis, Tsvetkov, Zettlemoyer, and Yih]{shi2023trusting}
Weijia Shi, Xiaochuang Han, Mike Lewis, Yulia Tsvetkov, Luke Zettlemoyer, and Scott Wen-tau Yih.
\newblock Trusting your evidence: Hallucinate less with context-aware decoding.
\newblock \emph{arXiv preprint arXiv:2305.14739}, 2023.

\bibitem[Shuster et~al.(2021)Shuster, Poff, Chen, Kiela, and Weston]{shuster-etal-2021-retrieval-augmentation}
Kurt Shuster, Spencer Poff, Moya Chen, Douwe Kiela, and Jason Weston.
\newblock Retrieval augmentation reduces hallucination in conversation.
\newblock In Marie-Francine Moens, Xuanjing Huang, Lucia Specia, and Scott Wen-tau Yih, editors, \emph{Findings of the Association for Computational Linguistics: EMNLP 2021}, pages 3784--3803, Punta Cana, Dominican Republic, November 2021. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2021.findings-emnlp.320}.
\newblock URL \url{https://aclanthology.org/2021.findings-emnlp.320}.

\bibitem[Team(2023)]{MosaicML2023Introducing}
MosaicML~NLP Team.
\newblock Introducing mpt-7b: A new standard for open-source, commercially usable llms, 2023.
\newblock URL \url{www.mosaicml.com/blog/mpt-7b}.
\newblock Accessed: 2023-05-05.

\bibitem[Tian et~al.(2023)Tian, Mitchell, Yao, Manning, and Finn]{tian2023fine}
Katherine Tian, Eric Mitchell, Huaxiu Yao, Christopher~D Manning, and Chelsea Finn.
\newblock Fine-tuning language models for factuality.
\newblock \emph{arXiv preprint arXiv:2311.08401}, 2023.

\bibitem[Touvron et~al.(2023{\natexlab{a}})Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozi{\`e}re, Goyal, Hambro, Azhar, et~al.]{touvron2023llama}
Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timoth{\'e}e Lacroix, Baptiste Rozi{\`e}re, Naman Goyal, Eric Hambro, Faisal Azhar, et~al.
\newblock Llama: Open and efficient foundation language models.
\newblock \emph{arXiv preprint arXiv:2302.13971}, 2023{\natexlab{a}}.

\bibitem[Touvron et~al.(2023{\natexlab{b}})Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et~al.]{touvron2023llamatwo}
Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et~al.
\newblock Llama 2: Open foundation and fine-tuned chat models.
\newblock \emph{arXiv preprint arXiv:2307.09288}, 2023{\natexlab{b}}.

\bibitem[van~der Poel et~al.(2022)van~der Poel, Cotterell, and Meister]{van2022mutual}
Liam van~der Poel, Ryan Cotterell, and Clara Meister.
\newblock Mutual information alleviates hallucinations in abstractive summarization.
\newblock \emph{arXiv preprint arXiv:2210.13210}, 2022.

\bibitem[Varshney et~al.(2023)Varshney, Yao, Zhang, Chen, and Yu]{varshney2023stitch}
Neeraj Varshney, Wenlin Yao, Hongming Zhang, Jianshu Chen, and Dong Yu.
\newblock A stitch in time saves nine: Detecting and mitigating hallucinations of llms by validating low-confidence generation, 2023.

\bibitem[Xu and Cohen(2023)]{xu2023lightweight}
Zhichao Xu and Daniel Cohen.
\newblock A lightweight constrained generation alternative for query-focused summarization.
\newblock In \emph{Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval}, pages 1745--1749, 2023.

\bibitem[Xu et~al.(2024)Xu, Lamba, Ai, Tetreault, and Jaimes]{xu2024counterfactual}
Zhichao Xu, Hemank Lamba, Qingyao Ai, Joel Tetreault, and Alex Jaimes.
\newblock Cfe2: Counterfactual editing for search result explanation.
\newblock In \emph{Proceedings of the 2024 ACM SIGIR International Conference on Theory of Information Retrieval}, pages 145--155, 2024.

\bibitem[Zhang et~al.(2019)Zhang, Kishore, Wu, Weinberger, and Artzi]{zhang2019bertscore}
Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian~Q Weinberger, and Yoav Artzi.
\newblock Bertscore: Evaluating text generation with bert.
\newblock \emph{arXiv preprint arXiv:1904.09675}, 2019.

\end{thebibliography}

\documentclass[table]{style/style}

\usepackage{microtype}
\usepackage{hyperref}
\usepackage{url}
\usepackage{booktabs}
\usepackage{graphicx}
\usepackage{lineno}
\usepackage{enumitem}
\usepackage{listings} %
\usepackage{svg}

\definecolor{ darkblue}{rgb}{0, 0, 0.5}
\hypersetup{colorlinks=true, citecolor=darkblue, linkcolor=darkblue, urlcolor=darkblue}

\usepackage{amssymb}
\usepackage{multirow}
\usepackage{bigdelim}
\usepackage{todonotes}
\usepackage{longtable}
\usepackage{tabularray}
\usepackage{wrapfig}
\usepackage[most]{tcolorbox} 
\usepackage{url}
\usepackage{xspace}
\usepackage{svg}
\usepackage[absolute]{textpos} %

\usepackage{fdsymbol}   %

\usepackage[utf8]{inputenc} %
\usepackage[T1]{fontenc}    %
\usepackage{url}            %
\usepackage{booktabs}       %
\usepackage{amsfonts}       %
\usepackage{nicefrac}       %
\usepackage{microtype}      %
\usepackage[table,dvipsnames]{xcolor}         %
\usepackage{amsmath}
\usepackage[most]{tcolorbox}
\usepackage{csquotes}
\usepackage{wrapfig}
\usepackage{siunitx}
\usepackage{graphicx}
\usepackage{arydshln}
\usepackage{wrapfig}
\usepackage{enumitem}
\usepackage{soul} %

\usepackage{multirow}
\usepackage{xspace}
\usepackage{adjustbox}
\usepackage{pifont}
\usepackage{caption}
\usepackage{makecell}
\usepackage{subcaption}
\usepackage{bold-extra}
\usepackage{url}
\usepackage{float}
\usepackage{pgf-pie}

\usepackage{hyperref}
\crefname{figure}{Figure}{Figures}
\crefname{section}{Section}{Sections}
\crefname{equation}{Equation}{Equations}
\crefname{appendix}{Appendix}{Appendice}
\crefname{table}{Table}{Tables}

\definecolor{linkcolor}{RGB}{0, 0, 128}
\hypersetup{
     colorlinks   = true,
     citecolor    = linkcolor,
     linkcolor    = linkcolor,
     urlcolor     = linkcolor,
}
\usepackage{pifont}%
\usepackage{listings}

\setlist[itemize]{leftmargin=*,itemsep=0em,parsep=0.3em,topsep=0.3em}

\DeclareUnicodeCharacter{2212}{\ensuremath{-}}

\usepackage{tikz}

\usepackage{setspace}

\usepackage{nicematrix}
\newcolumntype{L}[1]{>{\raggedright\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\newcolumntype{C}[1]{>{\centering\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\newcolumntype{R}[1]{>{\raggedleft\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\newcolumntype{P}[1]{>{\centering\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\colorlet{lightgray}{White!30!lightgray}
\colorlet{lightblue}{White!70!MidnightBlue}

\title{\LARGE Context-aware Decoding Reduces Hallucination in Query-focused Summarization}

\authorOne{Zhichao Xu}
\affiliation{University of Utah}
\contribution[]{Corresponds to: \texttt{zhichao.xu@utah.edu}}

\abstract{
Query-focused summarization (QFS) aims to provide a summary of a single document/multi documents that can satisfy the information needs of a given query. It is useful for various real-world applications, such as abstractive snippet generation or more recent retrieval augmented generation (RAG). A prototypical QFS pipeline consists of a retriever (sparse or dense retrieval) and a generator (usually a large language model). However, applying large language models (LLM) potentially leads to hallucinations, especially when the evidence contradicts the prior belief of LLMs. There has been growing interest in developing new decoding methods to improve generation quality and reduce hallucination. In this work, we conduct a large-scale reproducibility study on one recently proposed decoding method\,---\,Context-aware Decoding (CAD).
In addition to replicating CAD's experiments on news summarization datasets, we include experiments on QFS datasets, and conduct more rigorous analysis on computational complexity and hyperparameter sensitivity. Experiments with eight different language models show that performance-wise, CAD improves QFS quality by (1) reducing factuality errors/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores, while also at a cost of increased inference-time FLOPs and reduced decoding speed. The \href{https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs}{code implementation} based on Huggingface Library is made available
}

\begin{document}

\maketitle

\section{Introduction}
Query-focused summarization (QFS) aims to provide a summary of a single document/multiple documents that can satisfy the information need of a given query~\citep{xu2023lightweight,marchionini2006exploratory}. 
It is useful for various real-world applications, such as abstractive snippet generation~\citep{chen2020abstractive} or more recent retrieval augmented generation (RAG)~\citep{lewis2020retrieval,gao2024retrievalaugmentedgenerationlargelanguage}. 
A prototypical QFS pipeline consists of a retriever to retrieve the documents to use as evidence, and a generator (usually based on large language models) to read the documents and generate natural language answers/summaries.

While the idea of combining large language models (LLM) and information retrieval seems tempting~\citep{metzler2021rethinking}, a few problems persist. 
For example, deploying generative models requires more inference FLOPs and specialized hardwares (GPUs or TPUs); and more importantly, the usage of LLMs in QFS/RAG pipeline also potentially leads to the hallucination problem where the generated summary contains information contradicting the source documents/evidence~\citep{shah2022situating,bender2021dangers}, which might hurt users' experience in information seeking processes and potentially escalate the spread of misinformation. 
As such, mainstream search engines (Google, Bing, etc.) still deploy extractive snippets~\citep{bast2014efficient}.

In parallel to training better language models to reduce hallucinations, there has been growing interest in developing decoding methods to allow the language models to better attend to evidence or to penalize degeneration, thus improving generation quality. For example, \citet{li-etal-2023-contrastive} propose contrastive decoding, leveraging the output probability distribution difference between an expert LM and an amateur LM to improve generation quality. 
DExperts~\citep{liu-etal-2021-dexperts} focus on reducing toxicity by contrasting the expert LM with toxicity LM.
\cite{xu2023lightweight} show that directly injecting knowledge constraints can potentially produce more faithful summarization.
A few works~\citep{shi2023trusting,van2022mutual} make use of pointwise mutual information to let the LM focus more on the input context to reduce hallucination. 

In this work, we focus on Context-aware Decoding~\citep[CAD,][]{shi2023trusting}. 
CAD leverages the idea of pointwise mutual information and proposes a product-of-experts enhancement to the vanilla decoding method to make the generation more conditioned on the input evidence. 
We conduct a large-scale reproducibility study with different choices of language models to study CAD's effectiveness in query-focused summarization, and include more rigorous analysis on computational complexity and hyperparameter sensitivity.
Extensive experiments show that CAD improves QFS quality by (1) reducing factual mistakes/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores. 
On the other hand, CAD also introduces additional inference-time FLOPs and potentially slows down decoding speed.

\section{Background}
\label{section:background}
\paragraph{Context-aware Decoding.}
Consider a language model parameterized by $\theta$, a context $\mathbf{c}$, an input query $\mathbf{x}$, and we want to generate a piece of text $\mathbf{y}$. 
In QFS paradigm, $\mathbf{c}$ can be the source document, $\mathbf{x}$ and $\mathbf{y}$ will be the query and summary, respectively.
In the vanilla decoding scenario, each token of generation $\mathbf{y}$ is sampled autoregressively from the output probability distribution from language model $\theta$ conditioned on input query $\mathbf{x}$ and context $\mathbf{c}$:
\begin{equation}
    y_t \sim p_{\theta} (y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) \propto \text{Softmax}\big( \, \text{logit}_{\theta} (y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) / \tau \big)
\end{equation}
Here $\tau$ is a temperature hyperparameter to control the sharpness of the probability distribution over the candidate vocabulary $\mathcal{V}$.
Examples of the sampling strategy include greedy and Top-$p$ sampling~\citep{holtzman2019curious}.

As we are operating on a conditional generation problem, we hope the generation is more conditioned on input context $\mathbf{c}$. 
\cite{shi2023trusting} utilize the idea of pointwise mutual information (PMI).
We consider two events $\mathbf{c}$ and $y_t$:
\begin{equation}
    \mathbf{PMI} ( p_{\theta} (y_t ; \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) ) = \log \Big( \frac{ p_{\theta}(y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t})} { p_{\theta}(y_t | \mathbf{x}, \mathbf{y}_{<t})} \Big)
\end{equation}
The interpretation of the above equation is we would like to measure the association of event $y_t$ (predicting a specific token) and event $\mathbf{c}$ (the presence of context $\mathbf{c}$).
$p_{\theta}(y_t | \mathbf{x}, \mathbf{y}_{<t})$ is the prior probability, representing the model's prior belief from its parameters $\theta$ when no presence of context $\mathbf{c}$, whereas $p_{\theta}(y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t})$ is the posterior probability with $\mathbf{c}$.
To leverage PMI, \citet{shi2023trusting} propose to multiply PMI with the original probability distribution:
\begin{equation}
    y_t \sim \tilde{p}_{\theta} (y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) \propto p_{\theta} (y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) \Big( \frac{ p_{\theta}(y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t})} { p_{\theta}(y_t | \mathbf{x}, \mathbf{y}_{<t})} \Big)^{\alpha}
\end{equation}
Here $\alpha$ is a hyperparameter to control the PMI weighting.
We still need to normalize the above equation to acquire a valid probability distribution, rearranging the terms yields
\begin{equation}
    y_t \sim \text{Softmax} \big[ \big( (1+\alpha) \, \text{logit}_{\theta}(y_t | \mathbf{c}, \mathbf{x}, \mathbf{y}_{<t}) - \alpha \,\text{logit}_{\theta}(y_t | \mathbf{x}, \mathbf{y}_{<t}) \big) / \tau \big]
\end{equation}
Here larger $\alpha$ yields more weight on the context $\mathbf{c}$ and $\alpha=0$ reduces to vanilla decoding. 
\cite{shi2023trusting} recommend setting $\alpha=0.5$ for summarization tasks and we also examine this choice in Section~\ref{section:experiments}.

\paragraph{Analysis of Computational Cost.}
We follow the notations used in~\cite{kaplan2020scaling}. Let us consider the Transformer architecture with following hyperparameters: number of layers $n_{\text{layer}}$, dimension of the embedding layer and intermediate representation $d_{\text{model}}$, dimension of the intermediate feedforward layer $d_{\text{ff}}$, dimension of the attention output $d_{\text{attn}}$ and number of attention heads $n_{\text{heads}}$; and denote the number of tokens in the input sequence as $n_{\text{input}}$.
The model size $N$ is the number of \emph{non-embedding} parameters of the model, $N\approx |\theta|$:
\begin{equation}
    N \approx 2\, d_{\text{model}}\, n_{\text{layer}} (2\, d_{\text{attn}}+d_{\text{ff}}) = 12\, n_{\text{layer}} \,d_{\text{model}}^2  \,\,\,\, \text{with} \,\, d_{\text{attn}}=d_{\text{ff}}/4 = d_{\text{model}}
\end{equation}
The total FLOPs per token in one forward pass is approximately $C_{\text{forward}}=2N+2 n_{\text{layer}} \, n_{\text{input}} \, d_{\text{attn}}$ (see Section 2.1 in~\cite{kaplan2020scaling} for details). 
In LLMs, $d_{\text{model}} \gg n_{\text{input}}/12$, we can roughly ignore the second term and use $C_{\text{forward}}\approx 2N$.

With the above analysis, we know that performing one forward pass of CAD yields in total $(2\,|\mathbf{y}_{<t}|+2\,|\mathbf{x}|+|\mathbf{c}|) \cdot C_\text{forward}$ FLOPs compared to $(|\mathbf{y}_{<t}|+|\mathbf{x}| +|\mathbf{c}|) \cdot C_\text{forward}$ from vanilla decoding.

\section{Experiments}
\label{section:experiments}

We carry out experiments on two QFS datasets and two news summarization datasets, with different choices of backbone language models, including pre-trained language models and instruction finetuned language models.

\subsection{Prompting Templates}
For QFS task, we treat the source document as context $\mathbf{c}$, and treat the original query as query $\mathbf{x}$ and prompt the model to generate summarization $\mathbf{y}$. 
For the summarization task, the source document is used as context $\mathbf{c}$, and we manually append a query $\mathbf{x}$ in the shape of \texttt{[summarize the following document:]}. The detailed prompting templates can be found in Appendix~\ref{subsec:templates}.

\subsection{Experiment Setup}
\paragraph{Datasets.} We benchmark CAD's performance on two widely adopted QFS datasets. Debatepedia dataset is collected by~\cite{nema-etal-2017-diversity} from 663 debates of 53 diverse categories in an encyclopedia of debates and consists of 12K/0.7K/1.0K query-document summarization triplets $(\emph{query}, \emph{document}, \emph{summarization})$.
PubMedQA~\citep{jin-etal-2019-pubmedqa} is a long-form abstractive question-answering dataset from the biomedical domain with the contexts available.
We also use two news summarization datasets from the original CAD paper: CNN Dailymail~\citep{see-etal-2017-get} and XSUM~\citep{narayan-etal-2018-dont} for reproducibility purposes.

\paragraph{Language Models.} 
We study language models of different architecture choices (encoder-decoder models vs decoder-only models) and training paradigms (pre-trained vs instruction finetuned).
\begin{itemize}
    \item \texttt{LLaMA-7B}~\citep{touvron2023llama} is a decoder-only language model pre-trained with causal language modeling objective from FAIR.
    \item \texttt{Llama-2-7B}~\citep{touvron2023llamatwo} is the second iteration of \texttt{LLaMA} model by better pre-training corpus and minor architectural improvements. 
    \item \texttt{MPT-7B}~\citep{MosaicML2023Introducing} is a decoder-only language model pre-trained with causal language modeling objective from MosaicML
    \item \texttt{MPT-7B-instruct} is based on \texttt{MPT-7B} and further instruction finetuned with an instruction finetuning dataset \texttt{mosaicml/instruct-v1}.
    \item \texttt{Mistral-7B-v0.1}~\citep{jiang2023mistral} is a decoder-only language model pre-trained with causal language modeling objective from Mistral AI.
    \item \texttt{Mistral-7B-instruct-v0.1} is based on \texttt{Mistral-7B-v0.1} and further instruction finetuned with an internal dataset.
    \item \texttt{Flan-T5-xl}~\citep{chung2022scaling} (3B) is an encoder-decoder model from T5 family~\cite{raffel2020exploring}; it is pre-trained with span-infilling objective and further adapted with the causal language modeling objective. It is instruction finetuned on the Flan collection.
    \item \texttt{Flan-T5-xxl}~\citep{chung2022scaling} (11B) undergoes same training procedure as \texttt{Flan-T5-xl} and demonstrates strong performance on natural language understanding tasks. 
\end{itemize}
We acquire weights of all listed models from Huggingface library\footnote{\url{https://huggingface.co/models}}.

\paragraph{Evaluation.}
Similar to~\cite{shi2023trusting}, we evaluate the generation quality with ROUGE F1~\citep{lin-2004-rouge}, BERTScore-precision~\citep{zhang2019bertscore} and FactKB~\citep{feng2023factkb}. 
ROUGE measures the match of lexical patterns while BERTScore measures the semantic similarity between the generation and the reference.
FactKB is a model-based evaluation metric that classifies whether the generation is faithful to the document.
For BERTScore we use \texttt{roberta-large-mnli} as it shows a higher correlation with human judgments. 
For FactKB we use the original classifier trained by the authors~\citep{feng2023factkb}.

\paragraph{Hyperparameters}
Our experimental settings follow~\cite{shi2023trusting} to a large extent. 
For decoding, we set temperature=1, use a combination of Top-$k$ and Top-$p$ sampling, and vary $\alpha$ to study the effectiveness of CAD. 
For the main results reported in Section~\ref{sec:results}, we use $\alpha=0.3$ and use 0.15, 0.5 as an ablation study.
Detailed choice of hyperparameters can be found in Appendix~\ref{subsec:hyperparameters}.

\section{Results and Analysis}
\label{sec:results}

\begin{table*}[!t]
\centering
\caption{Vanilla decoding vs CAD with news summarization datasets. }
\resizebox{0.9\textwidth}{!}{
\begin{tabular}{llcrrrrr}
\toprule
\textbf{Datasets}  & \textbf{Model} & \textbf{Decoding} & \textbf{ROUGE-1} & \textbf{ROUGE-2} & \textbf{ROUGE-L} & \textbf{BERTScore-P} & \textbf{FactKB}\\ 
\hline
\multirow{16}{*}{CNN Dailymail} & \multirow{2}{*}{\texttt{LLaMA-7B}} & Vanilla & 27.6 & 8.7 & 24.5 & 73.0 & 85.8 \\
& & CAD & 30.4 & 10.8 & 27.1 & 73.5 & 90.9 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Llama-2-7B}} & Vanilla & 26.8 & 8.6 & 23.9 & 72.3 & 80.6 \\
& & CAD & 29.3 & 10.3 & 26.2 & 73.1 & 84.9 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B}} & Vanilla & 28.9 & 9.8 & 25.7 & 73.0 & 86.2 \\
& & CAD & 30.5 & 11.1 & 27.0 & 73.8 & 90.4 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B-Instruct}} & Vanilla & 31.0 & 10.6 & 27.4 & 74.2 & 89.2 \\
& & CAD & 32.2 & 11.9 & 28.5 & 74.9 & 94.0 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B}} & Vanilla & 28.1 & 9.3 & 24.9 & 72.8 & 78.5 \\
& & CAD & 29.3 & 10.6 & 25.9 & 73.7 & 84.0 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B-Instruct}} & Vanilla & 34.2 & 11.4 & 30.5 & 75.4 & 93.1 \\
& & CAD & 34.6 & 13.3 & 30.9 & 75.5 & 95.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xl}} & Vanilla & 29.0 & 10.0 & 25.6 & 75.6 & 86.0 \\
& & CAD & 30.8 & 11.6 & 27.5 & 75.9 & 85.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xxl}} & Vanilla & 27.0 & 8.4 & 24.0 & 74.5 & 82.3 \\
& & CAD & 28.8 & 10.0 & 25.8 & 75.1 & 81.8 \\
\midrule
\multirow{16}{*}{XSUM} & \multirow{2}{*}{\texttt{LLaMA-7B}} & Vanilla & 19.7 & 3.9 & 15.5 & 71.5 & 38.9 \\
& & CAD & 19.9 & 3.8 & 15.6 & 71.8 & 53.8 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Llama-2-7B}} & Vanilla & 19.5 & 3.9 & 15.3 & 71.5 & 47.2 \\
& & CAD & 20.0 & 4.0 & 15.6 & 71.7 & 58.1 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B}} & Vanilla & 19.5 & 3.7 & 15.2 & 71.6 & 56.2 \\
& & CAD & 20.0 & 3.8 & 15.5 & 72.3 & 63.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B-Instruct}} & Vanilla & 22.9 & 5.3 & 17.7 & 73.4 & 51.5 \\
& & CAD & 22.3 & 5.1 & 17.2 & 76.1 & 62.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B}} & Vanilla & 23.2 & 5.9 & 18.0 & 72.6 & 43.0 \\
& & CAD & 22.5 & 5.9 & 17.6 & 73.3 & 51.8 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B-Instruct}} & Vanilla & 25.8 & 6.7 & 20.0 & 74.2 & 55.7 \\
& & CAD & 25.2 & 6.5 & 19.5 & 73.9 & 63.4 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xl}} & Vanilla & 34.8 & 12.8 & 27.5 & 80.4 & 53.3 \\
& & CAD & 35.8 & 13.8 & 28.5 & 80.8 & 55.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xxl}} & Vanilla & 36.0 & 13.9 & 28.2 & 80.8 & 54.8 \\
& & CAD & 36.7 & 14.7 & 29.3 & 81.1 & 54.9 \\

\bottomrule
\end{tabular}
}
\label{tab:summarization_main}
\end{table*}

\begin{table*}[!t]
\centering
\caption{Vanilla decoding vs CAD with QFS datasets.}
\resizebox{0.9\textwidth}{!}{
\begin{tabular}{llcrrrrr}
\toprule
\textbf{Datasets}  & \textbf{Model} & \textbf{Decoding} & \textbf{ROUGE-1} & \textbf{ROUGE-2} & \textbf{ROUGE-L} & \textbf{BERTScore-P} & \textbf{FactKB}\\ 
\hline
\multirow{16}{*}{Dbpedia} & \multirow{2}{*}{\texttt{LLaMA-7B}} & Vanilla & 13.3 & 2.9 & 11.8 & 68.8 & 29.6 \\
& & CAD & 13.4 & 3.1 & 12.0 & 68.9 & 36.1 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Llama-2-7B}} & Vanilla & 12.0 & 2.5 & 10.8 & 68.6 & 27.3 \\
& & CAD & 12.7 & 2.7 & 11.2 & 68.4 & 34.0 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B}} & Vanilla & 13.1 & 2.7 & 11.6 & 70.0 & 33.4 \\
& & CAD & 13.3 & 3.0 & 12.0 & 70.4 & 37.9 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B-Instruct}} & Vanilla & 14.8 & 3.7 & 13.1 & 71.5 & 28.1 \\
& & CAD & 14.5 & 3.4 & 12.8 & 72.2 & 31.3 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B}} & Vanilla & 13.8 & 3.2 & 12.2 & 69.4 & 28.8 \\
& & CAD & 13.8 & 3.0 & 12.3 & 69.1 & 36.9 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B-Instruct}} & Vanilla & 18.2 & 5.3 & 15.9 & 71.6 & 42.6 \\
& & CAD & 17.1 & 4.7 & 15.1 & 71.1 & 41.6 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xl}} & Vanilla & 18.0 & 5.4 & 16.2 & 74.4 & 37.5 \\
& & CAD & 16.2 & 4.5 & 14.5 & 73.7 & 46.3 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xxl}} & Vanilla & 20.6 & 6.8 & 18.6 & 75.0 & 33.9 \\
& & CAD & 16.2 & 4.4 & 14.5 & 73.5 & 43.0 \\
\midrule
\multirow{16}{*}{PubMedQA} & \multirow{2}{*}{\texttt{LLaMA-7B}} & Vanilla & 25.5 & 7.6 & 21.3 & 71.2 & 79.5 \\
& & CAD & 26.0 & 8.2 & 21.6 & 71.3 & 86.4 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Llama-2-7B}} & Vanilla & 27.7 & 9.0 & 23.0 & 72.0 & 87.4 \\
& & CAD & 28.6 & 9.5 & 23.6 & 72.1 & 90.8 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B}} & Vanilla & 27.8 & 8.7 & 23.0 & 72.0 & 92.4 \\
& & CAD & 27.4 & 8.5 & 22.8 & 71.7 & 96.3 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{MPT-7B-Instruct}} & Vanilla & 29.0 & 9.7 & 24.0 & 72.9 & 96.4 \\
& & CAD & 27.6 & 8.7 & 22.6 & 72.2 & 97.3 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B}} & Vanilla & 25.0 & 8.0 & 20.8 & 71.7 & 67.9 \\
& & CAD & 24.4 & 7.6 & 20.1 & 71.5 & 75.7 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Mistral-7B-Instruct}} & Vanilla & 31.7 & 11.9 & 26.4 & 74.3 & 97.5 \\
& & CAD & 31.2 & 11.6 & 26.0 & 73.9 & 97.5 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xl}} & Vanilla & 32.3 & 11.4 & 26.6 & 76.1 & 95.7 \\
& & CAD & 30.5 & 10.1 & 24.9 & 75.3 & 95.2 \\ \cline{2-8}
& \multirow{2}{*}{\texttt{Flan-T5-xxl}} & Vanilla & 33.1 & 11.6 & 27.0 & 76.3 & 95.6 \\
& & CAD & 31.3 & 10.6 & 25.6 & 75.6 & 94.3 \\

\bottomrule
\end{tabular}
}
\label{tab:qfs_main}
\end{table*}

We try to study the following research questions:
\begin{itemize}
    \item RQ1: Performance-wise, can CAD (1) improve news summarization performance as discussed by~\cite{shi2023trusting}? and (2) reduce hallucination/factuality errors in QFS?
    \item RQ2: What is a good choice of hyperparameter $\alpha$?
    \item RQ3: What is the computational trade-off?
\end{itemize}
\paragraph{Quality of Generation.}
We present the results on QFS datasets and news summarization datasets in Table~\ref{tab:qfs_main} and Table~\ref{tab:summarization_main}, respectively.
On news summarization datasets, CAD achieves improved ROUGE scores with all eight models on CNN Dailymail dataset, and improves with six models on XSUM datasets. 
In terms of factuality errors, CAD improves FactKB scores on six out of eight models on CNN Dailymail dataset and eight out of eight models on XSUM dataset. 
These results support the finding in~\cite{shi2023trusting} that CAD can improve news summarization performance.

In terms of results on QFS datasets, we notice that while CAD can reduce hallucination/factuality errors as measured by FactKB score, the ROUGE score is not improved on Dbpedia dataset and slightly drops on PubMedQA dataset.
We hypothesize that the different performance patterns might be because the references in two QFS datasets are highly abstractive, and myopically conditioning on the input context may not achieve an improved ROUGE score.

As for performance between different language models, a clear observation is that instruction finetuned models (\texttt{MPT-7B-Instruct} and \texttt{Mistral-7B-Instruct-v0.1}) outperform their pre-trained counterparts. 
On the relatively challenging PubMedQA dataset, \texttt{Mistral-7B-Instruct-v0.1} achieves comparable performance to \texttt{Flan-T5-xxl}, despite \texttt{Flan-T5-xxl} being trained on more academic style NLP tasks. 

\paragraph{Choice of hyperparameter $\alpha$.}
We show a comparison between different choices of hyperparameter $\alpha$ against model performance in Table~\ref{tab:hyperparameter}. 
From QFS results in Table~\ref{tab:hyperparameter} we notice that there is a trade-off between FactKB score and ROUGE score as $\alpha$ increases, whereas on the two news summarization datasets, the ROUGE score and FactKB score are positively correlated. 
We hypothesize that the different trends are due to the design choices in building the datasets.
On the news summarization datasets, setting $\alpha=0.3$ yields an overall better performance.

\begin{table*}[!t]
\centering
\captionsetup{width=0.8\columnwidth}
\caption{Performance comparison between difference choices of hyperparameter $\alpha$, recall that $\alpha=0$ reduces to vanilla decoding. The reported numbers are averaged over eight models.}
\resizebox{0.8\columnwidth}{!}{
\begin{tabular}{llrrrrr}
\toprule
\textbf{Datasets}  & $\boldsymbol{\alpha}$ \textbf{value} & \textbf{ROUGE-1} & \textbf{ROUGE-2} & \textbf{ROUGE-L} & \textbf{BERTScore-P} & \textbf{FactKB}\\ 
\hline
\multirow{4}{*}{Dbpedia} & $\alpha$=0.0 & 15.5 & 4.1 & 13.8 & 71.2 & 32.7 \\
& $\alpha$=0.15 & 14.9 & 3.8 & 13.3 & 71.1 & 35.1 \\
& $\alpha$=0.3 & 14.7 & 3.6 & 13.1 & 70.1 & 38.1 \\
& $\alpha$=0.5 & 14.5 & 3.6 & 13.0 & 70.9 & 38.7 \\
\midrule
\multirow{4}{*}{PubMedQA} & $\alpha$=0.0 & 29.1 & 9.7 & 24.0 & 73.3 & 89.1 \\
& $\alpha$=0.15 & 28.7 & 9.6 & 23.8 & 73.2 & 90.6 \\
& $\alpha$=0.3 & 28.4 & 9.4 & 23.4 & 73.0 & 91.7 \\
& $\alpha$=0.5 & 27.9 & 9.1 & 23.1 & 72.7 & 93.3 \\
\midrule
\multirow{4}{*}{CNN Dailymail} & $\alpha$=0.0 & 29.1 & 9.6 & 25.8 & 73.9 & 85.2 \\
& $\alpha$=0.15 & 30.4 & 10.9 & 27.1 & 74.3 & 87.2 \\
& $\alpha$=0.3 & 30.7 & 11.2 & 27.4 & 74.4 & 88.4 \\
& $\alpha$=0.5 & 30.7 & 11.2 & 27.3 & 74.4 & 87.7 \\
\midrule
\multirow{4}{*}{XSUM} & $\alpha$=0.0 & 25.2 & 7.0 & 19.7 & 74.5 & 50.1 \\
& $\alpha$=0.15 & 25.4 & 7.2 & 19.8 & 74.8 & 55.7 \\
& $\alpha$=0.3 & 25.3 & 7.2 & 19.9 & 74.9 & 58.0 \\
& $\alpha$=0.5 & 24.8 & 6.9 & 19.4 & 74.9 & 58.1 \\
\bottomrule
\end{tabular}
}
\label{tab:hyperparameter}
\end{table*}

\paragraph{Reporting Decoding Speed.}
Previously we have analyzed the additional complexity in terms of FLOPs of CAD. 
Now we take a look at the difference in decoding speed on a real-world dataset.
The experiment is conducted on a server with a single Nvidia RTX 3090 GPU, CUDA version 12.1, and PyTorch version 2.1.0.
We implement two forward passes, first with $p_{\theta}(y_t|\mathbf{c}, \, \mathbf{x}, \, \mathbf{y}_{<t})$, and second with $p_{\theta}(y_t|\mathbf{x},\, \mathbf{y}_{<t})$; another option is packing two computation into one batch which will increase CUDA memory usage.

On CNN Dailymail dataset, vanilla decoding with FP16 and batch size=1 leads to a decoding speed of 0.042 second/token, while the decoding speed of CAD with FP16 and batch size=1 is 0.077 second/token with two forward passes. 
This suggests while CAD can improve generation quality, it also slows down decoding speed or requires more CUDA memory.

\section{Related Work}
\label{sec:related}

\paragraph{Search as an Information-seeking Process.}
\cite{marchionini2006exploratory} categorize search into three two different information seeking processes\,---\,\emph{Information Lookup} and \emph{Exploratory Search}.
Information lookup refers to the user looking for simple facts or asking factoid questions, while exploratory search/search to learn often involves multiple iterations.
Different approaches have been explored to reduce users' cognitive loads in the information-seeking process. 
For example, Perplexity AI\footnote{\url{https://www.perplexity.ai/}} transfers web search queries into QA-style questions, and cites sources in the answers to improve verifiability.
Snippets~\citep{bast2014efficient} provide short previews for each webpage to allow users to locate relevant information without examining the webpage.
Counterfactual explanations~\citep{xu2024counterfactual} suggest counterfactual queries as explanations and demonstrate search sessions' effectiveness can be improved.

\paragraph{Hallucination in Natural Language Generation.}
\citet{ji2023survey} present a survey of hallucination in natural language generation. 
There are mainly two types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. 
Intrinsic hallucination refers to the generated output that contradicts the source content, whereas extrinsic hallucination refers to the generated output that cannot be verified from the source content. 
In retrieval augmented generation, we mainly focus on intrinsic hallucination.
Pre-trained language models (not instruction finetuned) are more prone to hallucination, as the gigantic pre-training corpus may contain untruthful texts and the teacher forcing training procedure lets the models pick up this untruthful information.
\citet{shuster-etal-2021-retrieval-augmentation} show that retrieval augmentation can reduce hallucinations in downstream tasks.
\citet{tian2023fine} show that language models can be trained to reduce hallucination via reinforcement learning techniques.
\citet{varshney2023stitch} propose to use lm logits to detect hallucination in the generation process and mitigate it by retrieving evidence and rewriting.

\paragraph{Decoding Methods to Improve Generation Quality.}
\citet{li-etal-2023-contrastive} propose Contrastive Decoding (CD) to improve generation quality. The idea is to contrast the stronger expert model's prediction with a smaller amateur model. 
A recent work~\citep{o2023contrastive} shows that CD can also improve the reasoning ability of the language models on multiple reasoning benchmark datasets.
Based on a similar idea, \citet{liu-etal-2021-dexperts} show that by contrasting to a toxic model (language model trained with toxicity contents), the toxicity of the expert language model's generations can be reduced.
A different flavor of controlled generation aims to let the model generate specific tokens/phrases. 
\citet{lu-etal-2021-neurologic} show that such a goal can be achieved by utilizing specific data structure and interpolation of lm logits.

\section{Conclusion and Limitations}
In this work we carry out a reproducibility study of a recently proposed decoding method\,---\,Context-aware Decoding (CAD). 
Experiments on query-focused summarization datasets and news summarization datasets show that CAD can improve QFS and news summarization quality by reducing factual mistakes/hallucinations while mostly retaining the match of lexical patterns, but also with additional inference-time FLOPs and reduced decoding speed.
Due to the limited bandwidth and resources, we only experiment with language models no larger than 11B (\texttt{Flan-T5-xxl}), larger language models may exhibit different performance patterns.

\bibliographystyle{plainnat}
\bibliography{references}

\section{Appendix}
\label{sec:appendix}

\subsection{Prompting Templates}
\label{subsec:templates}
We report the prompting templates in Table~\ref{tab:appendix_templates}.

\begin{table*}[!h]
\centering
\caption{Prompting Templates}
\vspace{-5pt}
\resizebox{\columnwidth}{!}{
\begin{tabular}{llll}
\toprule
\textbf{Datasets} & \textbf{Model} & \textbf{Decoding} & \textbf{Templates} \\ 
\midrule
\multirow{4}{*}{Dbpedia} & Decoder-only & Vanilla &
\begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{document\}.} \\ \texttt{According to the Document, the one-sentence answer to the Question is:} \end{tabular}\\
\cline{2-4}

 & Decoder-only & CAD &
\begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{None\}.} \\ \texttt{According to the Document, the one-sentence answer to the Question is:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & Vanilla & \begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{document\}.} \\ \texttt{According to the Document, the one-sentence answer to the Question is:} \end{tabular}\\
\cline{2-4}
& Encoder-Decoder & CAD & \begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{None\}.} \\ \texttt{According to the Document, the one-sentence answer to the Question is:} \end{tabular}\\
\midrule

\multirow{4}{*}{PubMedQA} & Decoder-only & Vanilla &
\begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{document\}.} \\ \texttt{According to the Document, the detailed answer to the Question is:} \end{tabular}\\
\cline{2-4}

& Decoder-only & CAD &
\begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{document\}.} \\ \texttt{According to the Document, the detailed answer to the Question is:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & Vanilla & \begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{document\}.} \\ \texttt{According to the Document, the detailed answer to the Question is:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & CAD & \begin{tabular}[c]{@{}l@{}l} \texttt{Question: \{query\}. Document: \{None\}.} \\ \texttt{According to the Document, the detailed answer to the Question is:} \end{tabular}\\
\midrule

\multirow{4}{*}{CNN Dailymail} & Decoder-only & Vanilla &
\begin{tabular}[c]{@{}l@{}l} \texttt{News article: \{document\}.} \\ \texttt{Summary of the above news article:} \end{tabular}\\
\cline{2-4}

& Decoder-only & CAD &
\begin{tabular}[c]{@{}l@{}l} \texttt{News article: \{None\}.} \\ \texttt{Summary of the above news article:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & Vanilla & \begin{tabular}[c]{@{}l@{}l} \texttt{Summarize the following article in one or two sentences.} \\ \texttt{\{document\}:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & CAD & \begin{tabular}[c]{@{}l@{}l} \texttt{Summarize the following article in one or two sentences.} \\ \texttt{\{None\}:} \end{tabular}\\
\midrule

\multirow{4}{*}{XSUM} & Decoder-only & Vanilla &
\begin{tabular}[c]{@{}l@{}l} \texttt{News article: \{document\}.} \\ \texttt{Summary of the above news article:} \end{tabular}\\
\cline{2-4}

 & Decoder-only & CAD &
\begin{tabular}[c]{@{}l@{}l} \texttt{News article: \{None\}.} \\ \texttt{Summary of the above news article:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & Vanilla & \begin{tabular}[c]{@{}l@{}l} \texttt{Summarize the following article in one or two sentences.} \\ \texttt{\{document\}:} \end{tabular}\\
\cline{2-4}

& Encoder-Decoder & CAD & \begin{tabular}[c]{@{}l@{}l} \texttt{Summarize the following article in one or two sentences.} \\ \texttt{\{None\}:} \end{tabular}\\

\bottomrule
\end{tabular}
}
\label{tab:appendix_templates}
\end{table*}

\subsection{Hyperparameters}
\label{subsec:hyperparameters}
We report the hyperparameters in Table~\ref{tab:appendix_hyperaprameters}.
\begin{table*}[!h]
\centering
\caption{Decoding Hyperparameters}
\vspace{-5pt}
\resizebox{\columnwidth}{!}{
\begin{tabular}{lrrrrrrr}
\toprule
\textbf{Dataset} & \textbf{Top-k} & \textbf{Top-p} & \textbf{num\_beams} & \textbf{repetition\_penalty} & \textbf{temperature} & \textbf{min\_new\_tokens} & \textbf{max\_new\_tokens} \\
\midrule
Dbpedia & 50 & 0.9 & 1 & 1.0 & 1.0 & 5 & 30 \\
PubMedQA & 50 & 0.9 & 1 & 1.0 & 1.0 & 40 & 100 \\
CNN Dailymail & 50 & 0.9 & 1 & 1.0 & 1.0 & 30 & 70 \\
XSUM & 50 & 0.9 & 1 & 1.0 & 1.0 & 20 & 50 \\

\bottomrule
\end{tabular}
}
\label{tab:appendix_hyperaprameters}
\end{table*}

\end{document}